# A Linear Quantum Coupler for Clean Bosonic Control

### v1

@@ Appendix C Effect of junction and DC flux asymmetries
$\begin{split}U_{\text{full}}/E_{L}=\frac{1}{2}&(\theta_{c}-2\phi_{\Delta}/3)^{2}\\ -\beta_{\Sigma}\cos(\phi_{d})&\left[\cos(\phi_{\Delta}/3)\cos(\theta_{c})-\sin(\phi_{\Delta}/3)\sin(\theta_{c})\right]\\ +\beta_{\Delta}\sin(\phi_{d})&\left[\cos(\phi_{\Delta}/3)\sin(\theta_{c})+\sin(\phi_{\Delta}/3)\cos(\theta_{c})\right]\end{split}$ (34)
@@
Figure 7: LINC Kerr in the presence of asymmetries a, Kerr of the LINC mode at $\phi_{\text{DC}}=\pi/2$, computed by direct diagonalization. Black contour lines correspond to $\|K_{\text{LINC}}\|=250,500,750,\text{and }1000$ KHz respectively. b, Operating point $\phi_{d}$ where the LINC is Kerr-free, as a function of asymmetry. With a minor change in the symmetric DC flux, the Kerr induced by small asymmetry can be nulled. Black contour lines correspond to flux values from $298\%\times\pi/$ 2 to $2102\%\times\pi/$ 2 in $0.5\%$ increments around the asymmetry-free DC operating point. c, third-order nonlinearity of the LINC dipole mode, $g_{30}$. With finite junction asymmetry, the self-Kerr and cross-Kerr-free points will differ. Contour lines correspond to $\|g_{30}\|=3,6,9,\text{and }12$ MHz respectively.
@@
$\begin{split}K_{\text{asym}}&\approx\frac{1}{2}\frac{\partial^{2}U}{\partial\theta_{c}^{4}}\theta_{zpf}^{4}\Bigg{|}_{\theta_{c}=\theta_{c}^{\text{min}},\phi_{d}=\pi/2}\\ &\approx E_{c}\beta_{\Delta}\sin\left(\phi_{\Delta}-\beta_{\Delta}\right)\end{split}$ (39)
@@
$K_{\text{asym}}\approx E_{c}\beta_{\Delta}\sin\left(\phi_{\Delta}-(1+\sqrt{2})\beta_{\Delta}\right).$ (40)
@@
$\begin{split}U/E_{L}=&\frac{1}{2}(\theta_{c}-2\phi_{\Delta}/3)^{2}\\ +\beta_{\Sigma}\sin(\phi_{AC})&\left[\cos(\phi_{\Delta}/3)\cos(\theta_{c})-\sin(\phi_{\Delta}/3)\sin(\theta_{c}))\right]\\ +\beta_{\Delta}\cos(\phi_{AC})&\left[\cos(\phi_{\Delta}/3)\sin(\theta_{c})+\sin(\phi_{\Delta}/3)\cos(\theta_{c}))\right].\end{split}$ (42)

### v2

@@ Appendix C Effect of junction and DC flux asymmetries
$\begin{split}U_{\text{full}}/E_{L}=\frac{1}{2}&(\hat{\theta}_{c}-2\phi_{\Delta}/3)^{2}\\ -\beta_{\Sigma}\cos(\phi_{d})&\left[\cos(\phi_{\Delta}/3)\cos(\hat{\theta}_{c})-\sin(\phi_{\Delta}/3)\sin(\hat{\theta}_{c})\right]\\ +\beta_{\Delta}\sin(\phi_{d})&\left[\cos(\phi_{\Delta}/3)\sin(\hat{\theta}_{c})+\sin(\phi_{\Delta}/3)\cos(\hat{\theta}_{c})\right]\end{split}$ (34)
@@
Figure 7: LINC Kerr in the presence of asymmetries a, Kerr of the LINC mode at $\phi_{\text{DC}}=\pi/2$, computed by direct diagonalization. Black contour lines correspond to $|\alpha_{L}|=250,500,750,\text{and }1000$ KHz respectively. b, Operating point $\phi_{d}$ where the LINC is Kerr-free, as a function of asymmetry. With a minor change in the symmetric DC flux, the Kerr induced by small asymmetry can be nulled. Black contour lines correspond to flux values from $298\%\times\pi/$ 2 to $2102\%\times\pi/$ 2 in $0.5\%$ increments around the asymmetry-free DC operating point. c, third-order nonlinearity of the LINC dipole mode, $g_{30}$. With finite junction asymmetry, the self-Kerr and cross-Kerr-free points will differ. Contour lines correspond to $|g_{30}|=3,6,9,\text{and }12$ MHz respectively.
@@
$\begin{split}\alpha_{L}^{\text{asym}}&\approx\frac{1}{2}\frac{\partial^{2}U}{\partial\theta_{c}^{4}}\theta_{zpf}^{4}\Bigg{|}_{\theta_{c}=\theta_{c}^{\text{min}},\phi_{d}=\pi/2}\\ &\approx E_{c}\beta_{\Delta}\sin\left(\phi_{\Delta}-\beta_{\Delta}\right)\end{split}$ (39)
@@
$\alpha_{L}^{\text{asym}}\approx E_{c}\beta_{\Delta}\sin\left(\phi_{\Delta}-(1+\sqrt{2})\beta_{\Delta}\right).$ (40)
@@
$\begin{split}U/E_{L}=&\frac{1}{2}(\hat{\theta}_{c}-2\phi_{\Delta}/3)^{2}\\ +\beta_{\Sigma}\sin(\phi_{AC})&\left[\cos(\phi_{\Delta}/3)\cos(\hat{\theta}_{c})-\sin(\phi_{\Delta}/3)\sin(\hat{\theta}_{c}))\right]\\ +\beta_{\Delta}\cos(\phi_{AC})&\left[\cos(\phi_{\Delta}/3)\sin(\hat{\theta}_{c})+\sin(\phi_{\Delta}/3)\cos(\hat{\theta}_{c}))\right].\end{split}$ (42)